OpenStudy (anonymous):

Suppose M is the midpoint of line FG. MG=7x-15, FG=33, x=? How do you solve this??

hartnn (hartnn):

since M is mid-point MG=MF=FG/2 so 7x-15=33/2 can u simplify this further?

OpenStudy (anonymous):

Thank you so much! I finished it

hartnn (hartnn):

your welcome :)